OpenStudy (anonymous):

Let f(x)=x^2+7x+7 and g(x)=x+9. Compute f∘g(x)

OpenStudy (anonymous):

fog(x)=f(g(x))=f(x+9) \[f(x+9)=(x+9)^2+7(x+9)+9=x^2+18x+81+7x+63+9=x^2+25x+153\]

OpenStudy (anonymous):

\[x^2+25x+153\]